### v1

# Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms

Abstract
[…]
pacs:
03.67.Mn, 03.65.Ud, 42.50.Dv, 42.50.-p
I I INTRODUCTION
Uncertainties in the simultaneous measurements of incompatible observables of quantum systems make quantum systems drastically different from classical ones. Uncertainty principle, first formulated by Heisenberg Heisenberg forms a central part of our understanding of quantum mechanics. Many works have been done on further developments and applications of uncertainty principle with the advent and progress of quantum information theory Robertson -Horodecki1 .
Von Neumann entropy, also called entanglement entropy, plays a crucial role in the quantification of quantum entanglement. Entropic manifestation of quantum entanglement forms a significant part of quantum information theory Horodecki1 .
Spin squeezing is an important experimental tool that have applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics Jian .
In this paper we consider a general entangled state of two two-level atoms. We show that though the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole, thus violating the entropic inequalities Horodecki1 , but the uncertainties of the subsystems are less than the system as a whole for certain range of the superposing constants of the quantum state. This is in contradiction with the consensus that more is the entropy, more is the uncertainty and more is the disorder. We also establish analytically a general relationship between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the system as a whole. Now, since the spin squeezing and spectroscopic squeezing are experimentally measurable, so, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
In section II we briefly discuss the two-level atoms and pseudo-spin operators. In section III we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems. We show that though the von Neumann entropy of the composite state is less than those of the subsystems, but, the uncertainties of the composite state are greater than those of the subsystems for certain range of the superposing constants of the quantum state. In section IV we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the composite state. In section V we present the conclusion.
I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS
An atom has many electronic energy levels, but when it interacts with an external monochromatic electromagnetic radiation, the atom makes a transition from one of its energy level to the other. In this case, we mainly concentrate on those two energy levels and hence the atom is called as a two-level atom. Now, if among the assembly of $N$ such two-level atoms, the $n$-th atom has the upper and lower energy levels, denoted as $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then, we can construct the pseudo-spin operators (with $\hbar=1$),
$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$ (1)
$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$ (2)
$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$ (3)
such that,
$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$ (4)
and two more relations with cyclic changes in $x$, $y$ and $z$ Itano . For the entire system of two two-level atoms, we have collective pseudo-spin operators,
$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$ (5)
where each term in the above summations is multiplied by the identity operators for all of the other atoms Itano , Sakurai .
The individual atomic operators satisfy
$\big{[}\hat{J}_{1_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $0,~{}\big{[}\hat{J}_{1_{x}},\hat{J}_{1_{y}}\big{]}=i\hat{J}_{1_{z}},$
$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...$ (6)
where the subscripts $1$ and $2$ are for atoms $1$ and $2$ respectively.
As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$ and $\hat{J}_{z}$ satisfy,
$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z}$ (7)
and two more relations with cyclic changes in $x$, $y$ and $z$.
The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively.
The use of these collective spin operators as discussed in Eqs. (5) are widely used in the problems of quantum optics Korbicz1 -Miyake2 .
I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
We, now, consider a system of two entangled two-level atoms, say atoms $a$ and $b$. A general pure state of this system in Schmidt decomposition is given as
$|\psi_{ab}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle,$ (8)
where $|\psi_{ab}\rangle$in $H_{a}\otimes H_{b}$, $c_{i}$ are the Schmidt coefficients and $\{|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle\}$ is a orthonormal product basis. The Schmidt coefficients $c_{i}$ satisfy
$\sum_{i=1}^{2}{c_{i}}^{2}=1.$ (9)
For the system of two two-level atoms we take the orthonormal product basis, made up of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, as
$\Big{\{}\big{(}c_{3}|\frac{1}{2}\rangle+c_{4}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{5}|\frac{1}{2}\rangle+c_{6}|-\frac{1}{2}\rangle\big{)}\Big{\}}$
and
$\Big{\{}\big{(}c_{4}|\frac{1}{2}\rangle-c_{3}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{6}|\frac{1}{2}\rangle-c_{5}|-\frac{1}{2}\rangle\big{)}\Big{\}},$
where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively and for the sake of simplicity $c_{3}$, $c_{4}$, $c_{5}$ and $c_{6}$ are assumed to be real. Since the basis vectors are normalized, we have
${c_{3}}^{2}+{c_{4}}^{2}$ $=$ $1,$ (10)
${c_{5}}^{2}+{c_{6}}^{2}$ $=$ $1$ (11)
Therefore, the complete form of $|\psi_{ab}\rangle$ is
$|\psi_{ab}\rangle=c_{1}\bigg{[}\big{(}c_{3}\Big{|}\frac{1}{2}\Big{\rangle}+c_{4}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\otimes$
$\big{(}c_{5}\Big{|}\frac{1}{2}\Big{\rangle}+c_{6}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}+$
$c_{2}\bigg{[}\big{(}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}$
$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}$ (12)
The above quantum state can be written as
$|\psi_{ab}\rangle=\Big{(}c_{1}c_{3}c_{5}+c_{2}c_{4}c_{6}\Big{)}\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}$ (13)
For making the calculations simple we let
$\Big{(}c_{1}c_{3}c_{5}+c_{2}c_{4}c_{6}\Big{)}$ $=$ $A,$ (14)
$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}$ $=$ $B$ (15)
$\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}$ $=$ $C,$ (16)
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}$ $=$ $D,$ (17)
where
$A^{2}+B^{2}+C^{2}+D^{2}=1.$ (18)
In order to study the spin squeezing and spectroscopic squeezing properties of the system we need to calculate the coordinate frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$ (19)
$\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ are three mutually orthogonal unit vectors. The expectation values as mentioned above are to be calculated over the quantum state in Eq. (12).
Now, to understand it concretely we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ along the positive $x$, $y$ and $z$ axes respectively so that $\langle\hat{\mathbf{J}}\rangle$ is
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$ (20)
and hence the uncertainties are the standard deviations
$\Delta J_{x,y}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x,y}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x,y}|\psi_{ab}\rangle^{2}}.$ (21)
Now, since for an arbitrary quantum state, as that given in Eq. (12) the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi_{ab}\rangle$. These are
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x^{\prime},y^{\prime}}|\psi_{ab}\rangle^{2}}.$ (22)
Now, since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ after rotation of the coordinate frame points along the $z^{\prime}$ axis, we have
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle=0,$ (23)
and hence, the standard deviations reduce to
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle}.$ (24)
We now proceed to calculate $\Delta J_{x^{\prime},y^{\prime}}$. We find that
$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C)$ (25)
$\langle\psi_{ab}|\hat{J}_{y}|\psi_{ab}\rangle$ $=$ $0,$ (26)
$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2}.$ (27)
Therefore, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction we perform the following rotation.
$\hat{J}_{x^{\prime}}$ $=$ $\hat{J}_{x}\cos\theta-\hat{J}_{z}\sin\theta$ (28)
$\hat{J}_{y^{\prime}}$ $=$ $\hat{J}_{y}$ (29)
$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$ (30)
where,
$\cos\theta$ $=$ $\frac{\langle\hat{J}_{z}\rangle}{|\langle\hat{\mathbf{J}}\rangle|}.$ (31)
It can be verified using Eqs. (26), (28), (29) and (31) that
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ 00 0 (32)
$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$ (33)
and hence the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now along the $z^{\prime}$-axis.
Now, using Eqs. (28) and (31), we get
$\langle\hat{J}_{x^{\prime}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}\rangle|^{2}}\Big{[}\langle\hat{J}_{x}^{2}\rangle\langle\hat{J}_{z}\rangle^{2}+\langle\hat{J}_{z}^{2}\rangle\langle\hat{J}_{x}\rangle^{2}$
$-\langle\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\rangle\langle\hat{J}_{x}\rangle\langle\hat{J}_{z}\rangle\Big{]}.$ (34)
We find that for the quantum state in Eq. (12),
$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)},$ (35)
$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=A^{2}+D^{2},$ (36)
$\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle=(A-D)(B+C).$
Therefore, using Eqs. (25), (32), (34) to (I.2), we get
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}$ (38)
So, using Eqs. (24) and (38), we get
$\Delta{J}_{x^{\prime}}^{2}=\frac{1}{2}\big{(}A+D\big{)}^{2}.$ (39)
Similarly,
$\langle\psi_{ab}|\hat{J}_{y}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$ (40)
Therefore, using Eqs. (24), (26), (29), (33) and (40), we get
$\Delta{J}_{y^{\prime}}^{2}$ $=$ $\langle\psi_{ab}|\hat{J}_{y^{\prime}}^{2}|\psi_{ab}\rangle$ (41)
$=$ $\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$
Now, using Eqs. (14), (15), (16), (17), (39) and (41), we get
$\Delta{J}_{x^{\prime}}=\sqrt{\Big{(}\frac{1}{2}+c_{1}c_{2}\Big{)}}\Big{(}c_{3}c_{5}+c_{4}c_{6}\Big{)},$ (42)
$\Delta{J}_{y^{\prime}}=\sqrt{\Big{(}\frac{1}{2}-c_{1}c_{2}\Big{)}}.$ (43)
These are the coordinate frame independent inherent uncertainties of the system.
We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$ respectively.
Now the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
$\hat{\rho}_{a}=c_{1}^{2}\Big{[}c_{3}\Big{|}\frac{1}{2}\Big{\rangle}+c_{4}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{3}+$
$\Big{\langle}-\frac{1}{2}\Big{|}c_{4}\Big{]}+c_{2}^{2}\Big{[}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}$
$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}$ (44)
We find that
$\langle\hat{J}_{a_{x}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{x}})=c_{3}c_{4}(c_{1}^{2}-c_{2}^{2}),$ (45)
$\langle\hat{J}_{a_{y}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{y}})=0,$ (46)
$\langle\hat{J}_{a_{z}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{z}})=\frac{1}{2}(c_{1}^{2}-c_{2}^{2})(c_{3}^{2}-c_{4}^{2}).$
So, the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ lies in the $x-z$ plane. Therefore, we perform the rotation
$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1}$ (48)
$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}}$ (49)
$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$ (50)
where
$\cos\theta_{1}$ $=$ $\frac{\langle\hat{J}_{a_{z}}\rangle}{|\langle\hat{\mathbf{J}}_{a}\rangle|}.$ (51)
It can be checked that
$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ 00 0 (52)
$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$ (53)
and hence the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
$\Delta J_{a_{x^{\prime},y^{\prime}}}=\sqrt{\langle{\hat{J}_{a_{x^{\prime},y^{\prime}}}}^{2}\rangle}.$ (54)
Similar to Eq. (34), we have
$\langle\hat{J}_{a_{x^{\prime}}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}_{a}\rangle|^{2}}\Big{[}\langle\hat{J}_{a_{x}}^{2}\rangle\langle\hat{J}_{a_{z}}\rangle^{2}+\langle\hat{J}_{a_{z}}^{2}\rangle\langle\hat{J}_{a_{x}}\rangle^{2}$
$-\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle\langle\hat{J}_{a_{x}}\rangle\langle\hat{J}_{a_{z}}\rangle\Big{]}.$ (55)
We find that
$\langle\hat{J}_{a_{x}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{x}}^{2})=\frac{1}{4},$ (56)
$\langle\hat{J}_{a_{y}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{y}}^{2})=\frac{1}{4},$ (57)
$\langle\hat{J}_{a_{z}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{z}}^{2})=\frac{1}{4},$ (58)
$\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle=tr\{\hat{\rho}_{a}(\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}})\}$
$=0,$ (59)
Therefore, using Eqs. (45) to (I.2) and (54) to (59), we get
$\Delta J_{a_{x^{\prime}}}$ $=$ $\frac{1}{2},$ (60)
$\Delta J_{a_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (61)
Similarly, it can be shown that for the atom $b$ the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
$\Delta J_{b_{x^{\prime}}}$ $=$ $\frac{1}{2},$ (62)
$\Delta J_{b_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (63)
Now, the von Neumann entropy of the composite state in Eq. (12) is zero as it is a pure state. Therefore,
$S_{a,b}=0.$ (64)
But, the von Neumann entropies of the atoms $a$ and $b$, which are entangled are
$S_{a}=S_{b}=-c_{1}^{2}log_{2}c_{1}^{2}-c_{2}^{2}log_{2}c_{2}^{2}.$ (65)
We now plot the uncertainties in Eqs. (42) and (43) and the von Neumann entropy in Eq. (65) with respect to the constant $c_{1}$ in FIG. 1.
Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the atoms $a$ and $b$ with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along horizontal axis and $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}$ and $S_{b}$ along the vertical axis. The thin line, thick line and the dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}=S_{b}$ respectively.
We observe that though the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00, that is that of the composite state, in the whole range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG. 1, but both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the range $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$. The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$ and $0.258\geq c_{2}\geq 0.001$. Therefore, both the uncertainties of the composite state in Eq. (12) are greater than those of the individual entangled atoms which have been represented in Eqs. (60) to (63). This is in contradiction with the prevailing consensus that more is the entropy, more is the uncertainty and more is the disorder.
Therefore, we see from FIG. 1 that there is violation of the entropic inequalities Horodecki1 , as in this case we have
$S_{a,b}<S_{a},$ (66)
$S_{a,b}<S_{b},$ (67)
but
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$ (68)
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$ (69)
and
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$ (70)
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$ (71)
in the ranges
$0.001\leq c_{1}\leq 0.258~{}~{}and~{}~{}0.966\leq c_{1}\leq 0.999,$
$0.999\geq c_{2}\geq 0.966~{}~{}and~{}~{}0.258\geq c_{2}\geq 0.001$
We have obtained this feature for the general state of two entangled two-level atoms.
We also observe from FIG. 1 that when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.
I.3 IV GENERALIZED RELATION BETWEEN VON NEUMANN ENTROPY AND SQUEEZING PARAMETERS IN TWO-LEVEL ATOMS
We establish generalized operational measures of the von Neumann entropies of the entangled atoms. We do this by establishing generalized relations between the von Neumann entropies of the entangled atoms and spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian and spectroscopic squeezing parameter used in the context of Ramsey spectroscopy, introduced by Wineland et. al Wineland2 ; Jian .
Using Eqs. (9) and (43), we get
$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}$ (72)
This is a quadratic equation in $c_{1}^{2}$ whose roots are
$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (73)
Therefore, using Eqs. (9) and (73), we get
$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (74)
So, using Eqs. (65), (73) and (74), we get
$S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$ (75)
$log_{2}\Big{[}\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}-$
$\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}$
or,
$S_{a}$ $=$ $S_{b}$
$=$ $-\Big{[}\frac{1}{2}\pm\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}.$
Here $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case we have $N=2$ and hence $j=1$. Now,
$\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}}=\xi_{s}$ (77)
is the spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian , where $\overrightarrow{n}_{\perp}$ is in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$.
In this case the $\overrightarrow{n}_{\perp}$ are the $x^{\prime}$ and $y^{\prime}$ directions. Now we have observed from FIG. 1 that we have squeezing in the $y^{\prime}$-quadrature. So, we take $\Delta J_{\overrightarrow{n}_{\perp}}=\Delta J_{y^{\prime}}$. Therefore, in this case
$\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}=\xi_{s_{y}},$ (78)
and hence we get with $j=1$,
$S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}.$
The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et. al Wineland2 ; Jian is
$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (80)
In this case since $N=2$, and we have squeezing in the $y^{\prime}$-quadrature, we have
$\xi_{R_{y}}=\sqrt{2}\frac{\Delta J_{y^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (81)
We write Eq. (75) as
$S_{a}$ $=$ $S_{b}$
$=$ $-\Big{[}\frac{1}{2}\pm\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}.$
The Eqs. (LABEL:3.57) and (LABEL:3.60) give generalized operational measures of von Neumann entropies of the entangled atoms.
Figure 2: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along the horizontal axis and $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ along the vertical axis. The thick line, dashed line and the dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ respectively.
We see from FIG. 2 that as the spin squeezing and spectroscopic squeezing of the composite state increases the von Neumann entropies of the entangled parts increase and when there is $100\%$ spin squeezing of the composite state the von Neumann entropies of the entangled parts attain their peak values. Therefore, more is the squeezing of the composite state more is the von Neumann entropy of the entangled parts. Then as the squeezing of the composite state decreases the von Neumann entropy of the entangled parts decrease. This feature, as we see, is for the general state of two two-level atoms given in Eq. (12).
I.4 V CONCLUSION
We have calculated analytically the uncertainties for a general state of two entangled two-level atoms both for the composite state and the entangled parts. We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state, but for certain range of the superposing constants of the quantum state, the uncertainties of the entangled parts are less than that of the composite state. This is not in accordance with the prevailing consensus that more is the entropy, more is the uncertainty. This feature, as we see, is for the general state of two two-level atoms.
We have also established analytically the generalized relationship between the von Neumann entropies of the entangled parts and the uncertainties, spin squeezing and spectroscopic squeezing parameters, used in the context of Ramsey spectroscopy, of the composite state. This provides a generalized operational measure of the von Neumann entropies of the entangled atoms.

### v2

@@
-Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms
+Uncertainty, von Neumann Entropy, and Squeezing in a Bipartite State of Two-Level Atoms
 Abstract
 […]
-pacs:
-03.67.Mn, 03.65.Ud, 42.50.Dv, 42.50.-p
+keywords: uncertainty, von Neumann entropy, two-level atom, spin squeezing, spectroscopic squeezing
 I I INTRODUCTION
-Uncertainties in the simultaneous measurements of incompatible observables of quantum systems make quantum systems drastically different from classical ones.
-Uncertainty principle, first formulated by Heisenberg Heisenberg forms a central part of our understanding of quantum mechanics.
-Many works have been done on further developments and applications of uncertainty principle with the advent and progress of quantum information theory Robertson -Horodecki1 .
+Uncertainties in the simultaneous measurements of incompatible observables of quantum systems, make quantum systems drastically different from classical ones.
+Uncertainty principle, first formulated by Heisenberg Heisenberg , forms a central part of our understanding of quantum mechanics.
+Many works have been done on further developments and applications of uncertainty principle, with the advent and progress of quantum information theory Robertson -Horodecki1 .
 Von Neumann entropy, also called entanglement entropy, plays a crucial role in the quantification of quantum entanglement.
 Entropic manifestation of quantum entanglement forms a significant part of quantum information theory Horodecki1 .
-Spin squeezing is an important experimental tool that have applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics Jian .
-In this paper we consider a general entangled state of two two-level atoms.
-We show that though the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole, thus violating the entropic inequalities Horodecki1 , but the uncertainties of the subsystems are less than the system as a whole for certain range of the superposing constants of the quantum state.
-This is in contradiction with the consensus that more is the entropy, more is the uncertainty and more is the disorder.
-We also establish analytically a general relationship between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the system as a whole.
-Now, since the spin squeezing and spectroscopic squeezing are experimentally measurable, so, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
-In section II we briefly discuss the two-level atoms and pseudo-spin operators.
-In section III we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems.
-We show that though the von Neumann entropy of the composite state is less than those of the subsystems, but, the uncertainties of the composite state are greater than those of the subsystems for certain range of the superposing constants of the quantum state.
-In section IV we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the composite state.
-In section V we present the conclusion.
-I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS
-An atom has many electronic energy levels, but when it interacts with an external monochromatic electromagnetic radiation, the atom makes a transition from one of its energy level to the other.
-In this case, we mainly concentrate on those two energy levels and hence the atom is called as a two-level atom.
-Now, if among the assembly of $N$ such two-level atoms, the $n$-th atom has the upper and lower energy levels, denoted as $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then, we can construct the pseudo-spin operators (with $\hbar=1$),
-$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$ (1)
-$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$ (2)
-$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$ (3)
+Spin squeezing is an important experimental tool that has potential applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics.
+Spin squeezing is the reduction of quantum fluctuations in collective spin observables of atomic systems, typically measured by the variance of the spin components.
+This is achieved by correlations between the spins, which can be generated through nonlinear interactions or entanglement Kitagawa ; Jian .
+Spectroscopic squeezing, on the other hand, refers to the reduction of quantum fluctuations in the spectral response of an atomic ensemble.
+This is typically measured by the variance of the absorption or emission spectrum.
+It has significant applications in enhancing resolutions and precisions in optical measurements, which lead to more accurate determination of atomic and molecular energy levels.
+Squeezing of spectroscopic signals enables the detection of weaker signals and more precise characterization of atomic and molecular systems.
+It enables the achievement of quantum-limited spectroscopy, where the measurement precision is limited only by the fundamental laws of quantum mechanics.
+It helps to develop more accurate atomic clocks, which rely on precise spectroscopic measurements.
+In quantum information processing, spectroscopic squeezing enables the development of more precise quantum gates and quantum computing architectures Wineland2 ; Jian .
+In this paper, we consider a general entangled state of two two-level atoms.
+We show that, although the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole —thus violating the entropic inequalities Horodecki1 —the uncertainties of the subsystems are lower than those of the system as a whole for certain ranges of the superposition constants of the quantum state.
+This is in contradiction with the idea that, the more is the entropy, more is the uncertainty, and the more is the disorder.
+We also establish, analytically, generalized relationships between the von Neumann entropies of the subsystems, and the uncertainties, spin squeezing, and spectroscopic squeezing parameters of the system as a whole.
+Now, since spin squeezing and spectroscopic squeezing are experimentally measurable, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
+It is worthwhile to mention here that noise and decoherence are significant practical challenges in experiments involving spin and spectroscopic squeezing.
+Noise and decoherence can destroy spin coherence, introduce spin fluctuations, and limit the squeezing factors in spin squeezing experiments.
+In spectroscopic squeezing experiments, they can broaden spectral lines, introduce spectral fluctuations, and limit squeezing factors.
+However, a lot of experimental and theoretical progress has been made in the study of spin and spectroscopic squeezing in many physical systems —for example, in ion traps and Bose —Einstein condensates [33 - 58].
+In section II, we briefly discuss two-level atoms, pseudo-spin operators, and spin and spectroscopic squeezing.
+In section III, we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems.
+We show that, although the von Neumann entropy of the composite state is less than those of the subsystems, the uncertainties of the composite state are greater than those of the subsystems for certain ranges of the superposing constants of the quantum state.
+In section IV, we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing, and spectroscopic squeezing parameters of the composite state.
+In section V, we present the conclusion.
+I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS, SPIN AND SPECTROSCOPIC SQUEEZING
+An atom has many electronic energy levels, but when it interacts with external monochromatic electromagnetic radiation, it undergoes a transition from one energy level to another.
+In this case, we mainly focus on those two energy levels; hence, the atom is referred to as a two-level atom.
+Now, if among an assembly of $N$ such two-level atoms, the $n$-th atom has upper and lower energy levels denoted by $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then we can construct the pseudo-spin operators (with $\hbar=1$),
+$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$
+$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$
+$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$
 such that,
-$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$ (4)
-and two more relations with cyclic changes in $x$, $y$ and $z$ Itano . For the entire system of two two-level atoms, we have collective pseudo-spin operators,
-$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$ (5)
-where each term in the above summations is multiplied by the identity operators for all of the other atoms Itano , Sakurai .
+$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$
+and two more relations with cyclic changes in $x$, $y$ and $z$ Itano .
+Since these operators are not actual spin angular momentum operators, but satisfy the same commutation relations as those of spin-$\frac{1}{2}$ particles, they are referred to as pseudo-spin operators.
+For the entire system of two two-level atoms, we define the collective pseudo-spin operators
+$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},~{}~{}and~{}~{}\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$
+where each term in the summations is implicitly tensor-multiplied by identity operators acting on the other atom Itano , Sakurai .
 The individual atomic operators satisfy
 $\big{[}\hat{J}_{1_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $0,~{}\big{[}\hat{J}_{1_{x}},\hat{J}_{1_{y}}\big{]}=i\hat{J}_{1_{z}},$
-$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...$ (6)
-where the subscripts $1$ and $2$ are for atoms $1$ and $2$ respectively.
-As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$ and $\hat{J}_{z}$ satisfy,
-$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z}$ (7)
-and two more relations with cyclic changes in $x$, $y$ and $z$.
-The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively.
-The use of these collective spin operators as discussed in Eqs. (5) are widely used in the problems of quantum optics Korbicz1 -Miyake2 .
+$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...,$
+where the subscripts $1$ and $2$ are for atoms $1$ and $2$, respectively.
+As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$, and $\hat{J}_{z}$ satisfy
+$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z},$ (1)
+and two more relations with cyclic changes in $x$, $y$, and $z$.
+The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$, respectively.
+The use of these collective spin operators, as discussed in Eq. (I.1), is widely used in problems of quantum optics Korbicz1 -Miyake2 .
+In order to study the spin squeezing and spectroscopic squeezing properties of a quantum state, say $|\psi\rangle$, of a system of two-level atoms, we need to calculate the coordinate-frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ along two mutually orthogonal directions in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
+$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$
+$\hat{n}_{1}$, $\hat{n}_{2}$, and $\hat{n}_{3}$ are three mutually orthogonal unit vectors.
+The expectation values mentioned above are to be calculated with respect to the quantum state $|\psi\rangle$.
+To understand this concretely, we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$, and $\hat{n}_{3}$ along the positive $x$, $y$, and $z$ axes, respectively, so that
+$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$
+and hence the uncertainties are given by the standard deviations
+$\Delta J_{x,y}=\sqrt{\langle\psi|{\hat{J}_{x,y}}^{2}|\psi\rangle-\langle\psi|\hat{J}_{x,y}|\psi\rangle^{2}}.$ (2)
+Now, since, for an arbitrary quantum state $|\psi\rangle$, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis, and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi\rangle$.
+Since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$, after rotation of the coordinate frame, points along the $z^{\prime}$ axis, we have
+$\langle\psi|\hat{J}_{x^{\prime}}|\psi\rangle=\langle\psi|\hat{J}_{y^{\prime}}|\psi\rangle=0,$
+and hence the standard deviations, as mentioned in Eq. (2), reduce to
+$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi\rangle}.$ (3)
+In this paper, we calculate these coordinate-independent uncertainties for a general pure state of two two-level atoms, both for the composite state and for the individual entangled parts.
+For a system of $N$ two-level atoms, the spin squeezing parameter, introduced by Kitagawa et al. Kitagawa ; Jian , is defined as
+$\xi_{s}=\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}},$
+where $\overrightarrow{n}_{\perp}$ lies in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$, and $j=\frac{N}{2}$.
+Now, conventionally, we take $\overrightarrow{n}_{\perp}$ as the directions of $x^{\prime}$ and $y^{\prime}$ axes, and hence, the spin squeezing parameters become
+$\xi_{s_{x}}=\sqrt{\frac{2}{j}}\Delta J_{x^{\prime}}~{}~{}and~{}~{}\xi_{s_{y}}=\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}.$ (4)
+If, for a quantum state, $\xi_{s_{x}}<1$, then we have $\xi_{s_{y}}>1$, respecting Heisenberg’s uncertainty principle, and we conclude that we have spin squeezing in the $x$-quadrature —and vice versa.
+The degree of spin squeezing is quantified by this spin squeezing parameter.
+The smaller $\xi_{s}$ is (i.e., the more $\xi_{s}<1$), the greater the spin squeezing.
+The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et al. Wineland2 ; Jian , is given by
+$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$
+In a similar fashion, as mentioned above, we take the directions of $\overrightarrow{n}_{\perp}$ as the directions of $x^{\prime}$ and $y^{\prime}$ axes, and hence, the spectroscopic squeezing parameters become
+$\xi_{R_{x}}=\sqrt{N}\frac{\Delta J_{x^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}~{}~{}and~{}~{}\xi_{R_{y}}=\sqrt{N}\frac{\Delta J_{y^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (5)
+If $\xi_{R_{x}}$ or $\xi_{R_{y}}$ $<1$, we have spectroscopic squeezing in the system.
+Spectroscopic squeezing is quantified using these parameters.
+The smaller the value of $\xi_{R}<1$, the greater the spectroscopic squeezing.
+We now proceed to study the coordinate-independent inherent uncertainties and the von Neumann entropy of a general system of two two-level atoms.
 I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
-We, now, consider a system of two entangled two-level atoms, say atoms $a$ and $b$. A general pure state of this system in Schmidt decomposition is given as
-$|\psi_{ab}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle,$ (8)
-where $|\psi_{ab}\rangle$in $H_{a}\otimes H_{b}$, $c_{i}$ are the Schmidt coefficients and $\{|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle\}$ is a orthonormal product basis.
+A general pure state of two systems, say $A$ and $B$, is given by
+$|\psi_{AB}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{A}\rangle\otimes|e^{i}_{B}\rangle,$
+where $|\psi_{AB}\rangle$in $H_{A}\otimes H_{B}$, $c_{i}$ are the Schmidt coefficients, and $\{|e^{i}_{A}\rangle\otimes|e^{i}_{B}\rangle\}$ form an orthonormal product basis.
 The Schmidt coefficients $c_{i}$ satisfy
 $\sum_{i=1}^{2}{c_{i}}^{2}=1.$ (6)
-For the system of two two-level atoms we take the orthonormal product basis, made up of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, as
+We now consider a system of two entangled two-level atoms, denoted as atoms $a$ and $b$.
+To construct the composite quantum pure state $|\psi_{ab}\rangle$ of the two atoms, we take the orthonormal product basis, composed of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, given by
 $\Big{\{}\big{(}c_{3}|\frac{1}{2}\rangle+c_{4}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{5}|\frac{1}{2}\rangle+c_{6}|-\frac{1}{2}\rangle\big{)}\Big{\}}$
@@
 $\Big{\{}\big{(}c_{4}|\frac{1}{2}\rangle-c_{3}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{6}|\frac{1}{2}\rangle-c_{5}|-\frac{1}{2}\rangle\big{)}\Big{\}},$
-where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively and for the sake of simplicity $c_{3}$, $c_{4}$, $c_{5}$ and $c_{6}$ are assumed to be real.
+where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$, corresponding to eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$, respectively.
+For the sake of simplicity, $c_{3}$, $c_{4}$, $c_{5}$, and $c_{6}$ are assumed to be real.
 Since the basis vectors are normalized, we have
-${c_{3}}^{2}+{c_{4}}^{2}$ $=$ $1,$ (10)
-${c_{5}}^{2}+{c_{6}}^{2}$ $=$ $1$ (11)
+${c_{3}}^{2}+{c_{4}}^{2}=1$
+and
+${c_{5}}^{2}+{c_{6}}^{2}=1.$
 Therefore, the complete form of $|\psi_{ab}\rangle$ is
@@
 $c_{2}\bigg{[}\big{(}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}$
-$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}$ (12)
+$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}.$
 The above quantum state can be written as
@@
 $\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
-$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}$ (13)
-For making the calculations simple we let
+$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}.$ (7)
+To simplify the calculations, we let
 $\Big{(}c_{1}c_{3}c_{5}+c_{2}c_{4}c_{6}\Big{)}$ $=$ $A,$ (8)
-$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}$ $=$ $B$ (15)
+$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}$ $=$ $B,$ (9)
 $\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}$ $=$ $C,$ (10)
@@
 $A^{2}+B^{2}+C^{2}+D^{2}=1.$ (12)
-In order to study the spin squeezing and spectroscopic squeezing properties of the system we need to calculate the coordinate frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
-$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$ (19)
-$\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ are three mutually orthogonal unit vectors. The expectation values as mentioned above are to be calculated over the quantum state in Eq. (12).
-Now, to understand it concretely we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ along the positive $x$, $y$ and $z$ axes respectively so that $\langle\hat{\mathbf{J}}\rangle$ is
-$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$ (20)
-and hence the uncertainties are the standard deviations
-$\Delta J_{x,y}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x,y}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x,y}|\psi_{ab}\rangle^{2}}.$ (21)
-Now, since for an arbitrary quantum state, as that given in Eq.
-(12) the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi_{ab}\rangle$.
-These are
-$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x^{\prime},y^{\prime}}|\psi_{ab}\rangle^{2}}.$ (22)
-Now, since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ after rotation of the coordinate frame points along the $z^{\prime}$ axis, we have
-$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle=0,$ (23)
-and hence, the standard deviations reduce to
-$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle}.$ (24)
-We now proceed to calculate $\Delta J_{x^{\prime},y^{\prime}}$. We find that
-$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C)$ (25)
+Therefore, Eq. (7) becomes
+$|\psi_{ab}\rangle=A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+$
+$C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}.$ (13)
+We now proceed to calculate the coordinate independent inherent uncertainties, $\Delta J_{x^{\prime},y^{\prime}}$, as mentioned in the previous section.
+To that end, we first calculate $\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$, $\langle\psi_{ab}|\hat{J}_{y}|\psi_{ab}\rangle$, and $\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$.
+We find that
+$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C),$ (14)
 $\langle\psi_{ab}|\hat{J}_{y}|\psi_{ab}\rangle$ $=$ $0,$ (15)
-$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2}.$ (27)
-Therefore, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction we perform the following rotation.
+$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2},$ (16)
+and hence, the square of the magnitude of the mean spin vector is
+$|\langle\hat{\mathbf{J}}\rangle|^{2}=\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle^{2}+\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle^{2}$
+$=A^{2}(A^{2}+B^{2}+C^{2}+2BC)+$
+$D^{2}(B^{2}+C^{2}+D^{2}-2A^{2}+2BC)+$
+$2AD(B+C)^{2}.$ (17)
+We notice from Eqs. (14) to (16) that the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction, we perform the following rotation:
 $\hat{J}_{x^{\prime}}$ $=$ $\hat{J}_{x}\cos\theta-\hat{J}_{z}\sin\theta$ (18)
 $\hat{J}_{y^{\prime}}$ $=$ $\hat{J}_{y}$ (19)
-$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$ (30)
-where,
+$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$
+where
 $\cos\theta$ $=$ $\frac{\langle\hat{J}_{z}\rangle}{|\langle\hat{\mathbf{J}}\rangle|}.$ (20)
-It can be verified using Eqs. (26), (28), (29) and (31) that
-$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ 00 0 (32)
-$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$ (33)
-and hence the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now along the $z^{\prime}$-axis.
-Now, using Eqs. (28) and (31), we get
+It can be verified using Eqs. (15), (18), (19), and (20) that
+$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$
+$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$
+and hence, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now aligned along the $z^{\prime}$-axis.
+Using Eqs. (18) and (20), we obtain
 $\langle\hat{J}_{x^{\prime}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}\rangle|^{2}}\Big{[}\langle\hat{J}_{x}^{2}\rangle\langle\hat{J}_{z}\rangle^{2}+\langle\hat{J}_{z}^{2}\rangle\langle\hat{J}_{x}\rangle^{2}$
 $-\langle\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\rangle\langle\hat{J}_{x}\rangle\langle\hat{J}_{z}\rangle\Big{]}.$ (21)
-We find that for the quantum state in Eq. (12),
-$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)},$ (35)
-$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=A^{2}+D^{2},$ (36)
-$\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle=(A-D)(B+C).$
-Therefore, using Eqs. (25), (32), (34) to (I.2), we get
-$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}$ (38)
-So, using Eqs. (24) and (38), we get
+We now calculate $\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle$, $\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle$, and $\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle$ for the quantum state in Eq.
+(I.2), using the expression for $|\psi_{ab}\rangle$ given in Eq.
+(13).
+Now,
+$\hat{J}_{x}=\frac{1}{2}\Big{(}\hat{J}_{+}+\hat{J}_{-}\Big{)}.$
+Therefore,
+$\hat{J}_{x}^{2}=\frac{1}{4}\Big{(}\hat{J}_{+}^{2}+\hat{J}_{-}^{2}+\hat{{J}_{+}}\hat{{J}_{-}}+\hat{{J}_{-}}\hat{{J}_{+}}\Big{)},$ (22)
+and
+$\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}=\frac{1}{2}\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}+$
+$\hat{J}_{-}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{-}\Big{)}.$ (23)
+Since we are dealing here only with two atoms, $a$ and $b$, we have
+$\hat{J}_{+}$ $=$ $\hat{J}_{a_{+}}+\hat{J}_{b_{+}},$
+$\hat{J}_{-}$ $=$ $\hat{J}_{a_{-}}+\hat{J}_{b_{-}},$
+$\hat{J}_{z}$ $=$ $\hat{J}_{a_{z}}+\hat{J}_{b_{z}}.$ (24)
+Therefore,
+$\hat{J}_{+}^{2}=\hat{J}_{a_{+}}^{2}+\hat{J}_{b_{+}}^{2}+2\hat{J}_{a_{+}}\hat{J}_{b_{+}},$ (25)
+$\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}=\hat{J}_{a_{+}}\hat{J}_{a_{-}}+\hat{J}_{a_{-}}\hat{J}_{a_{+}}$
+$+\hat{J}_{b_{+}}\hat{J}_{b_{-}}+\hat{J}_{b_{-}}\hat{J}_{b_{+}}+2\hat{J}_{a_{+}}\hat{J}_{b_{-}}+2\hat{J}_{a_{-}}\hat{J}_{b_{+}},$
+and
+$\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}=\hat{J}_{a_{+}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{+}}+$
+$\hat{J}_{b_{+}}\hat{J}_{b_{z}}+\hat{J}_{b_{z}}\hat{J}_{b_{+}}+2\Big{(}\hat{J}_{a_{+}}\hat{J}_{b_{z}}+\hat{J}_{a_{z}}\hat{J}_{b_{+}}\Big{)}.$
+Now, as $A$, $B$, $C$, and $D$ are real, we have, using Eqs. (13) and (25),
+$\langle\psi_{ab}|\hat{J}_{+}^{2}|\psi_{ab}\rangle=\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+$
+$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+$
+$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}\Big{(}\hat{J}_{a_{+}}^{2}+\hat{J}_{b_{+}}^{2}+$
+$2\hat{J}_{a_{+}}\hat{J}_{b_{+}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
+$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
+$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
+$=2AD.$ (28)
+So,
+$\langle\psi_{ab}|\hat{J}_{-}^{2}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{+}^{2}|\psi_{ab}\rangle^{*}$
+$=2AD.$ (29)
+Using Eqs. (13) and (26), we obtain
+$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=$
+$\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+$
+$\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}$
+$\Big{(}\hat{J}_{a_{+}}\hat{J}_{a_{-}}+\hat{J}_{a_{-}}\hat{J}_{a_{+}}+\hat{J}_{b_{+}}\hat{J}_{b_{-}}+\hat{J}_{b_{-}}\hat{J}_{b_{+}}+$
+$2\hat{J}_{a_{+}}\hat{J}_{b_{-}}+2\hat{J}_{a_{-}}\hat{J}_{b_{+}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
+$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
+$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
+$=2\Big{(}A^{2}+B^{2}+C^{2}+D^{2}+2BC\Big{)}.$ (30)
+Now, by virtue of Eq. (12), Eq. (30) reduces to
+$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=2\Big{(}1+2BC\Big{)}.$
+Therefore, using Eqs. (22), (28), (29), and (31), we get
+$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)}.$ (32)
+Now, using Eqs. (13) and (24), we obtain
+$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+$
+$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+$
+$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}\Big{(}\hat{J}_{a_{z}}^{2}+\hat{J}_{b_{z}}^{2}+$
+$2\hat{J}_{a_{z}}\hat{J}_{b_{z}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
+$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
+$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
+$=A^{2}+D^{2}.$ (33)
+Similarly, using Eqs. (13) and (27), we get
+$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=\big{(}A-D\big{)}\times$
+$\big{(}B+C\big{)}.$ (34)
+Therefore,
+$\langle\psi_{ab}|\Big{(}\hat{J}_{-}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{-}\Big{)}|\psi_{ab}\rangle=$
+$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle^{*}=$
+$\big{(}A-D\big{)}\big{(}B+C\big{)}.$ (35)
+So, using Eqs. (23), (34), and (35), we obtain
+$\langle\psi_{ab}|\Big{(}\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\Big{)}|\psi_{ab}\rangle=$
+$(A-D)(B+C).$ (36)
+Therefore, using the expressions for $\langle\hat{J}_{x}\rangle$, $\langle\hat{J}_{z}\rangle$, $|\langle\hat{\mathbf{J}}\rangle|^{2}$, $\langle\hat{J}_{x}^{2}\rangle$, $\langle\hat{J}_{z}^{2}\rangle$, and $\langle\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\rangle$ from Eqs.
+(14), (16), (17), (32), (33), and (36), respectively, in Eq.
+(21), and simplifying, we get
+$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}.$ (37)
+So, using Eqs. (3) and (37), we have
 $\Delta{J}_{x^{\prime}}^{2}=\frac{1}{2}\big{(}A+D\big{)}^{2}.$ (38)
-Similarly,
+Now,
+$\hat{J}_{y}=\frac{1}{2i}\Big{(}\hat{J}_{+}-\hat{J}_{-}\Big{)}.$
+Therefore,
+$\langle\psi_{ab}|\hat{J}_{y}^{2}|\psi_{ab}\rangle=\frac{1}{4}\langle\psi_{ab}|\Big{(}\hat{{J}_{+}}\hat{{J}_{-}}+\hat{{J}_{-}}\hat{{J}_{+}}$
+$-\hat{J}_{+}^{2}-\hat{J}_{-}^{2}\Big{)}|\psi_{ab}\rangle.$ (39)
+So, using Eqs. (28), (29), and (31) in Eq. (39), and simplifying, we obtain
 $\langle\psi_{ab}|\hat{J}_{y}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$ (40)
-Therefore, using Eqs. (24), (26), (29), (33) and (40), we get
+Therefore, using Eqs. (3), (19), and (40), we get
 $\Delta{J}_{y^{\prime}}^{2}$ $=$ $\langle\psi_{ab}|\hat{J}_{y^{\prime}}^{2}|\psi_{ab}\rangle$ (41)
 $=$ $\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$
-Now, using Eqs. (14), (15), (16), (17), (39) and (41), we get
+Now, using Eqs. (8), (9), (10), (11), (38), and (41), we get
 $\Delta{J}_{x^{\prime}}=\sqrt{\Big{(}\frac{1}{2}+c_{1}c_{2}\Big{)}}\Big{(}c_{3}c_{5}+c_{4}c_{6}\Big{)},$ (42)
 $\Delta{J}_{y^{\prime}}=\sqrt{\Big{(}\frac{1}{2}-c_{1}c_{2}\Big{)}}.$ (43)
-These are the coordinate frame independent inherent uncertainties of the system.
-We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$ respectively.
-Now the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
+These are the coordinate frame-independent inherent uncertainties of the composite state $|\psi_{ab}\rangle$ of the system.
+We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$, and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$, respectively.
+Now, the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
 $\hat{\rho}_{a}=c_{1}^{2}\Big{[}c_{3}\Big{|}\frac{1}{2}\Big{\rangle}+c_{4}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{3}+$
 $\Big{\langle}-\frac{1}{2}\Big{|}c_{4}\Big{]}+c_{2}^{2}\Big{[}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}$
-$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}$ (44)
+$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}.$
 We find that
@@
 So, the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ lies in the $x-z$ plane. Therefore, we perform the rotation
-$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1}$ (48)
-$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}}$ (49)
-$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$ (50)
+$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1},$ (47)
+$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}},$ (48)
+$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$
 where
 $\cos\theta_{1}$ $=$ $\frac{\langle\hat{J}_{a_{z}}\rangle}{|\langle\hat{\mathbf{J}}_{a}\rangle|}.$ (49)
-It can be checked that
-$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ 00 0 (52)
-$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$ (53)
-and hence the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
+It can be verified, using Eqs. (45), (47), (48), and (49), that
+$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ $0,$
+$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$
+and hence, the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
 $\Delta J_{a_{x^{\prime},y^{\prime}}}=\sqrt{\langle{\hat{J}_{a_{x^{\prime},y^{\prime}}}}^{2}\rangle}.$ (50)
-Similar to Eq. (34), we have
+Similar to Eq. (21), we have, using Eqs. (47) and (49),
 $\langle\hat{J}_{a_{x^{\prime}}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}_{a}\rangle|^{2}}\Big{[}\langle\hat{J}_{a_{x}}^{2}\rangle\langle\hat{J}_{a_{z}}\rangle^{2}+\langle\hat{J}_{a_{z}}^{2}\rangle\langle\hat{J}_{a_{x}}\rangle^{2}$
@@
 $\langle\hat{J}_{a_{z}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{z}}^{2})=\frac{1}{4},$ (54)
-$\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle=tr\{\hat{\rho}_{a}(\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}})\}$
-$=0,$ (59)
+$\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle=$
+$tr\{\hat{\rho}_{a}(\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}})\}=0.$ (55)
 Therefore, using Eqs. (44) to (I.2) and (50) to (55), we get
@@
 $\Delta J_{a_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (57)
-Similarly, it can be shown that for the atom $b$ the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
+Similarly, it can be shown that, for atom $b$, the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
 $\Delta J_{b_{x^{\prime}}}$ $=$ $\frac{1}{2},$ (58)
 $\Delta J_{b_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (59)
-Now, the von Neumann entropy of the composite state in Eq. (12) is zero as it is a pure state. Therefore,
-$S_{a,b}=0.$ (64)
-But, the von Neumann entropies of the atoms $a$ and $b$, which are entangled are
+Now, the von Neumann entropy of the composite state in Eq. (I.2) is zero, as it is a pure state. Therefore,
+$S_{a,b}=0.$
+But the von Neumann entropies of atoms $a$ and $b$, which are entangled, are
 $S_{a}=S_{b}=-c_{1}^{2}log_{2}c_{1}^{2}-c_{2}^{2}log_{2}c_{2}^{2}.$ (60)
-We now plot the uncertainties in Eqs. (42) and (43) and the von Neumann entropy in Eq. (65) with respect to the constant $c_{1}$ in FIG. 1.
-Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the atoms $a$ and $b$ with respect to the constant $c_{1}$.
-Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$.
-$c_{1}$ has been plotted along horizontal axis and $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}$ and $S_{b}$ along the vertical axis.
-The thin line, thick line and the dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}=S_{b}$ respectively.
-We observe that though the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00, that is that of the composite state, in the whole range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG.
-1, but both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the range $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$.
-The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$ and $0.258\geq c_{2}\geq 0.001$.
+We now plot the uncertainties in Eqs. (42) and (43), and the von Neumann entropy in Eq. (60), with respect to the constant $c_{1}$ in FIG. 1.
+Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of atoms $a$ and $b$ with respect to the constant $c_{1}$.
+Here, we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$.
+The constant $c_{1}$ is plotted along the horizontal axis, while $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the entropies $S_{a}$ and $S_{b}$ are plotted along the vertical axis.
+The thin line, thick line, and dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and $S_{a}=S_{b}$, respectively.
+We observe that, although the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00 —that is, greater than that of the composite state —over the entire range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG.
+1, both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the ranges $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$.
+The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$, and $0.259\geq c_{2}\geq 0.045$, respectively.
 Therefore, both the uncertainties of the composite state in Eq.
-(12) are greater than those of the individual entangled atoms which have been represented in Eqs.
+(I.2) are greater than those of the individual entangled atoms, which have been represented in Eqs.
 (56) to (59).
-This is in contradiction with the prevailing consensus that more is the entropy, more is the uncertainty and more is the disorder.
+This is in contradiction with the prevailing idea that the more the entropy, the greater the uncertainty and disorder.
 Therefore, we see from FIG. 1 that there is violation of the entropic inequalities Horodecki1 , as in this case we have
-$S_{a,b}<S_{a},$ (66)
-$S_{a,b}<S_{b},$ (67)
+$S_{a,b}<S_{a},$
+$S_{a,b}<S_{b},$
 but
-$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$ (68)
-$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$ (69)
+$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$
+$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$
 and
-$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$ (70)
-$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$ (71)
+$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$
+$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$
 in the ranges
 $0.001\leq c_{1}\leq 0.258~{}~{}and~{}~{}0.966\leq c_{1}\leq 0.999,$
-$0.999\geq c_{2}\geq 0.966~{}~{}and~{}~{}0.258\geq c_{2}\geq 0.001$
+$0.999\geq c_{2}\geq 0.966~{}~{}and~{}~{}0.259\geq c_{2}\geq 0.045$
 We have obtained this feature for the general state of two entangled two-level atoms.
 We also observe from FIG.
-1 that when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.
+1 that, when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value, while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.
+Figure 2: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of atoms $a$ and $b$ with respect to the constant $c_{1}$.
+Here, we have taken $c_{3}=\frac{3}{8}$, $c_{4}=\frac{\sqrt{55}}{8}$, $c_{5}=\frac{4}{7}$ and $c_{6}=\frac{\sqrt{33}}{7}$.
+The constant $c_{1}$ is plotted along the horizontal axis, while $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the entropies $S_{a}$ and $S_{b}$ are plotted along the vertical axis.
+The thick line, dashed line, and thin line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and $S_{a}=S_{b}$, respectively.
+In FIG.
+1, we have chosen the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ as $\frac{1}{\sqrt{2}}$; that is, the quantum state $|\psi_{ab}\rangle$ is symmetric with respect to these parameters.
+Now, it can be seen that the above-mentioned feature also holds for asymmetric combinations of these parameters.
+This is because we observe from Eqs.
+(42) and (43) that while $\Delta J_{x^{\prime}}$ depends on the parameters $c_{1},c_{2},c_{3},c_{4},c_{5}$, and $c_{6}$, $\Delta J_{y^{\prime}}$ depends only on $c_{1}$ and $c_{2}$.
+Therefore, variations in the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ do not affect $\Delta J_{y^{\prime}}$, but they do affect $\Delta J_{x^{\prime}}$.
+Thus, $\Delta J_{y^{\prime}}$ is always greater than that of the individual entangled atoms in the ranges $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$.
+However, whether $\Delta J_{x^{\prime}}$ will be greater than that of the individual entangled atoms in the above-mentioned ranges of $c_{1}$ depends on the choice of the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$.
+Now, in Eq.
+(42), the minimum value of the factor $\sqrt{\Big{(}\frac{1}{2}+c_{1}c_{2}\Big{)}}$ is $0.707$.
+Therefore, $\Delta J_{x^{\prime}}$ is greater than $\frac{1}{2}$ —that is, greater than that of the individual enatngled atoms —as long as the factor $(c_{3}c_{5}+c_{4}c_{6})$ in Eq.
+(42) is greater than $0.707$.
+Thus, it does not matter whether the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ are symmetric or asymmetric, as long as their combination makes the factor $(c_{3}c_{5}+c_{4}c_{6})$ in Eq.
+(42) greater than $0.707$.
+As an example, in FIG.
+2, we plot $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ for an asymmetric combination of the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$, where $c_{3}=\frac{3}{8}$, $c_{4}=\frac{\sqrt{55}}{8}$, $c_{5}=\frac{4}{7}$, and $c_{6}=\frac{\sqrt{33}}{7}$.
+We observe the same result: that both $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$, for the composite state, are greater than those of the individual entangled atoms in the same ranges of the parameters $c_{1}$ and $c_{2}$, as mentioned previously.
 I.3 IV GENERALIZED RELATION BETWEEN VON NEUMANN ENTROPY AND SQUEEZING PARAMETERS IN TWO-LEVEL ATOMS
-We establish generalized operational measures of the von Neumann entropies of the entangled atoms.
-We do this by establishing generalized relations between the von Neumann entropies of the entangled atoms and spin squeezing parameter introduced by Kitagawa et.
-al Kitagawa ; Jian and spectroscopic squeezing parameter used in the context of Ramsey spectroscopy, introduced by Wineland et.
-al Wineland2 ; Jian .
-Using Eqs. (9) and (43), we get
-$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}$ (72)
-This is a quadratic equation in $c_{1}^{2}$ whose roots are
-$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (73)
-Therefore, using Eqs. (9) and (73), we get
-$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (74)
-So, using Eqs. (65), (73) and (74), we get
+We now establish generalized relationships between the von Neumann entropies of the entangled subsystems of two-level atoms and the spin and spectroscopic squeezing parameters of the composite state.
+This provides operational measures for the von Neumann entropies of the entangled subsystems of two-level atoms.
+Using Eqs. (6) and (43), we obtain
+$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}.$
+This is a quadratic equation in $c_{1}^{2}$, whose roots are
+$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}.$ (61)
+Therefore, using Eqs. (6) and (61), we obtain
+$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}.$ (62)
+Thus, using Eqs. (60), (61), and (62), we obtain
 $S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$ (63)
@@
 $\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$
-$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}$
-or,
+$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]},$
+or equivalently,
 $S_{a}$ $=$ $S_{b}$
@@
 $log_{2}\Big{[}\frac{1}{2}\mp\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}.$
-Here $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case we have $N=2$ and hence $j=1$. Now,
-$\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}}=\xi_{s}$ (77)
-is the spin squeezing parameter introduced by Kitagawa et.
-al Kitagawa ; Jian , where $\overrightarrow{n}_{\perp}$ is in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$.
-In this case the $\overrightarrow{n}_{\perp}$ are the $x^{\prime}$ and $y^{\prime}$ directions.
-Now we have observed from FIG.
-1 that we have squeezing in the $y^{\prime}$-quadrature.
-So, we take $\Delta J_{\overrightarrow{n}_{\perp}}=\Delta J_{y^{\prime}}$.
-Therefore, in this case
-$\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}=\xi_{s_{y}},$ (78)
-and hence we get with $j=1$,
+Here, $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case, we have $N=2$, and hence $j=1$.
+Now, using Eq. (4) in Eq. (LABEL:3.55), and taking $j=1$, we obtain
 $S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}\times$
@@
 $log_{2}\Big{[}\frac{1}{2}\mp\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}.$
-The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et. al Wineland2 ; Jian is
-$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (80)
-In this case since $N=2$, and we have squeezing in the $y^{\prime}$-quadrature, we have
+This is the generalized relationship between the von Neumann entropies of the entangled subsystems of two-level atoms and the spin squeezing parameter.
+To construct a general relation between the von Neumann entropies of the entangled subsystems of two-level atoms and the spectroscopic squeezing parameter, we consider Eq.
+(5).
+In this case, since $N=2$, we have
 $\xi_{R_{y}}=\sqrt{2}\frac{\Delta J_{y^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (66)
-We write Eq. (75) as
+Using Eq. (66), we rewrite Eq. (63) as
 $S_{a}$ $=$ $S_{b}$
@@
 $log_{2}\Big{[}\frac{1}{2}\mp\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}.$
-The Eqs. (LABEL:3.57) and (LABEL:3.60) give generalized operational measures of von Neumann entropies of the entangled atoms.
-Figure 2: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$.
-Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$.
-$c_{1}$ has been plotted along the horizontal axis and $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ along the vertical axis.
-The thick line, dashed line and the dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ respectively.
+Thus, we have established a generalized relationship between the von Neumann entropies of the entangled subsystems of two-level atoms and the spectroscopic squeezing parameter.
+The Eqs.
+(LABEL:3.57) and (LABEL:3.60) provide generalized operational measures of the von Neumann entropies of the entangled atoms.
+In FIG.
+3, we plot the spin and spectroscopic squeezing parameters of the composite state, and the von Neumann entropies of the entangled subsystems of two-level atoms, with respect to the constant $c_{1}$.
+Figure 3: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$.
+Here, we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$.
+The constant $c_{1}$ is plotted along the horizontal axis, while $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$, and $S_{a}=S_{b}$ are plotted along the vertical axis.
+The thick line, dashed line, and dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$, and $S_{a}=S_{b}$, respectively.
 We see from FIG.
-2 that as the spin squeezing and spectroscopic squeezing of the composite state increases the von Neumann entropies of the entangled parts increase and when there is $100\%$ spin squeezing of the composite state the von Neumann entropies of the entangled parts attain their peak values.
-Therefore, more is the squeezing of the composite state more is the von Neumann entropy of the entangled parts.
-Then as the squeezing of the composite state decreases the von Neumann entropy of the entangled parts decrease.
-This feature, as we see, is for the general state of two two-level atoms given in Eq.
-(12).
+3 that as the spin squeezing and spectroscopic squeezing of the composite state increase, the von Neumann entropies of the entangled parts also increase, and when there is $100\%$ spin squeezing of the composite state, the von Neumann entropies of the entangled parts attain their peak values.
+Therefore, the greater the squeezing of the composite state, the larger the von Neumann entropies of the entangled parts.
+Then, as the squeezing of the composite state decreases, the von Neumann entropies of the entangled parts also decrease.
+This feature, as we can see, applies to the general state of two two-level atoms given in Eq.
+(I.2).
 I.4 V CONCLUSION
-We have calculated analytically the uncertainties for a general state of two entangled two-level atoms both for the composite state and the entangled parts.
-We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state, but for certain range of the superposing constants of the quantum state, the uncertainties of the entangled parts are less than that of the composite state.
-This is not in accordance with the prevailing consensus that more is the entropy, more is the uncertainty.
-This feature, as we see, is for the general state of two two-level atoms.
-We have also established analytically the generalized relationship between the von Neumann entropies of the entangled parts and the uncertainties, spin squeezing and spectroscopic squeezing parameters, used in the context of Ramsey spectroscopy, of the composite state.
-This provides a generalized operational measure of the von Neumann entropies of the entangled atoms.
+We have analytically calculated the uncertainties for a general state of two entangled two-level atoms, both for the composite state and the entangled parts.
+We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state.
+However, for certain ranges of the superposition constants of the quantum state, the uncertainties of the entangled parts are found to be less than those of the composite state.
+This observation does not align with the prevailing idea that greater entropy implies greater uncertainty.
+As we can see, this feature holds for the general state of two two-level atoms.
+We have also analytically established generalized relationships between the von Neumann entropies of the entangled parts of two-level atoms and the uncertainties, spin squeezing, and spectroscopic squeezing parameters —used in the context of Ramsey spectroscopy —of the composite state.
+This provides generalized operational measures of the von Neumann entropies of the entangled atoms.